# Benchmarking the Text-to-SQL Capability of Large Language Models: A Comprehensive Evaluation

### v1

@@ 4.2 SQL Debugging
In recent studies, researchers have demonstrated that LLMs possess self-reflection and self-correction capabilities similar to those of humans [53, 41, 57]. Additionally, previous studies also investigate the potential of LLMs to debug the code they generate [4, 33]. In this section, we provide a comprehensive analysis of the performance of numerous SQL debugging methods across LLMs.
@@ 4.2.2 Debug Evaluation
Dataset $\longrightarrow$ SQLCoder-34B InternLM-70B Codellama-34B Llama2-Chat-70B
@@
Regenerate 0.00 13.21 12.42 18.75
w/ Wrong SQL 0.66 6.92 2.48 7.95
w/ Wrong SQL + System_error_info 4.64 9.43 11.18 11.36
w/ Wrong SQL + All_error_info 4.64 8.81 11.80 10.80
@@
Regenerate 9.27 0.63 11.80 15.34
w/ Wrong SQL 7.95 5.66 10.56 9.66
w/ Wrong SQL + System_error_info 10.60 8.81 10.56 11.93
w/ Wrong SQL + All_error_info 11.92 8.81 11.80 13.07
@@ 4.3 SQL Optimization
Execution efficiency of SQL queries is a critical aspect, particularly in real-time systems that utilize large-scale databases. In this section, we further explore whether LLMs are able to enhance the execution efficiency of correct SQL queries. Formally, the SQL optimization process [31] involves transforming the initial SQL queries $\mathcal{Y}$ into an optimized form, denoted as $\mathcal{Y}^{\rm o}$, with the goal of improving efficiency while maintaining identical results:
@@
VES is commonly employed to evaluate the efficiency of SQL query execution. However, in practice, LLMs can sometimes rewrite a correct SQL query into an incorrect one, making it challenging to figure out if the main reason for the decline in VES is due to these incorrect rewrites or a decrease in the SQL execution efficiency. Therefore, we propose a complementary metric C-VES (Correct-VES):
@@
Methods Prompt Template Metrics SQLCoder-34B InternLM-70B Codellama-34B InternLM2-20B
Baseline SimpleDDL-MD-Chat VES 24.28 22.63 18.94 19.81
C-VES 101.17 110.39 99.68 94.33
Two-Stage Generation zero-shot with $\mathcal{Y}$ VES 16.86 15.33 17.34 20.28
C-VES 102.19 102.17 102.08 101.42
w/ $\mathcal{Y}+\mathcal{S}+\mathcal{Q}$ VES 18.90 18.44 19.28 20.77
C-VES 102.14 102.43 101.47 101.30
few-shot w/ demo VES 18.84 17.97 18.34 20.86
C-VES 101.84 102.68 101.88 101.75
w/ demo + comments VES 18.28 17.45 18.52 20.92
C-VES 101.54 102.66 102.86 102.06
Direct Generation SimpleDDL-MD-Chat-Efficiency VES 27.77 25.23 20.75 25.93
C-VES 102.84 102.98 103.75 101.70
@@ 4.4 SQL-to-Text
The goal of SQL-to-Text is to transform the SQL query back into its original natural language question [51, 30, 42]. While it seems that SQL-to-Text cannot serve as a sub-task within the Text-to-SQL pipeline to improve the performance of End-to-End Text-to-SQL systems, employing this conversion as a supplementary step within the pipeline can indeed provide valuable insights. By converting the generated SQL statements back into text and juxtaposing these with the semantics of the original user questions, we can assess the accuracy of the SQL statements produced. In addition, it can assist researchers in evaluating the semantic comprehension capabilities of different LLMs, thus facilitating the development of more effective Text-to-SQL methodologies.
@@
SQLCoder-34B Codellama-34B InternLM-70B InternLM2-20B Llama2-Chat-70B
Rouge-1 0.118 0.359 0.408 0.410 0.368
Rouge-2 0.039 0.167 0.198 0.294 0.159
Rouge-L 0.108 0.320 0.372 0.358 0.325
BertScore 0.779 0.856 0.886 0.887 0.885
LLM Evaluator 22.5% 49.8% 65.8% 71.3% 60%
To this end, we assess the performance of SQL-to-Text across different LLMs (See Appendix A.4 for prompt templates). The selected metrics for evaluation encompass the F1 values of Rouge-1/2/L and BertScore, along with the application of LLM to assess the semantic coherence between the two texts. The evaluation results are depicted in Table 7. InternLM2 demonstrates the highest performance, followed by InternLM, while Codellama and SQLCoder exhibit comparatively lower performance. This highlights that even in regards to semantic description of code, general-purpose models exhibit significantly stronger descriptive capabilities compared to coding-specific models.
@@ 4.5 Schema Linking
Schema linking is recognized as a crucial prerequisite of generating correct SQL queries. It involves aligning entity references in the question with the corresponding schema tables or columns, requiring the model to understand both structure and value of the database, as well as the the semantics of user questions. In LLM-based Text-to-SQL, prior studies [9, 33] design prompt instructions with in-context learning examples to enable LLMs to retrieve linked tables and columns, which are then used for the downstream Text-to-SQL task. However, none of these methods individually evaluate the performance of schema linking, and explicit evaluation metrics have yet to be established. Moreover, despite considerable advancements in semantic comprehension and generalization brought by LLMs, the performance of schema linking is still far from promising.
@@ 4.5.1 Evaluation Metric for Schema Linking: RES
What schema linking method is considered good? A straightforward goal of schema linking is that the GT tables should be retrieved as much as possible. However, due to the ambiguity inherent in natural language and the potential semantic similarities between candidate tables, retrieving more GT tables typically comes at the cost of a higher redundancy, known as the precision-recall trade-off. As discussed in [33], excessive table retrieval can introduce redundant joins between tables, potentially impairing the EX of Text-to-SQL generation.
@@
where $T_{n}$ and $\hat{T}_{n}$ denote the set of GT tables and retrieved tables for the $n$-th instance, respectively, $|\cdot|$ refers to the scale of a set.
@@ 4.5.2 Schema Linking Evaluation
  • Zero Shot: A schema linking prompt proposed in C3 [9], which instructs the LLM to rank all the tables from relevant to irrelevant in a zero-shot manner.
  • Few Shot: Instructing the LLM to retrieve only the most important tables with few-shot demonstration, presented in DIN-SQL [33].
@@
Note that we organize the information of the database schema in a way similar to "SimpleDDL-" in the prompt of all the methods mentioned above, which ignores the information about foreign keys. However, as argued in [46], foreign keys embody features of known schema relations and provide importance clues for understanding the database structure. To this end, we conduct experiments under both settings, w/ and w/o foreign keys, to investigate how incorporating foreign keys in the prompt influences the performance of schema linking. Results are demonstrated in Table 8 (refer to Appendix B for results on detailed metrics like Exact Match & Subset Match).
@@
Can foreign key information facilitate schema linking? The introduction of foreign key information yield improved performance across all methods and all LLMs. This is evident since a valid JOIN operation in SQL queries is typically based on foreign keys. The foreign key information helps the model retrieve more GT tables by indicating all potential table pairs involved in a JOIN operation.
@@ A.4 SQL-to-Text
Listing LABEL:lst:sql_to_text is the prompt template used for SQL-to-Text task. Listing LABEL:lst:Semantic_consistency is utilized to assess the consistency between the transformed text description and the original user query semantics. To mitigate the influence of the order of the two text inputs, we employ InternLM2 to conduct two sequential tests and calculate the average of the results.
@@ A.5 Schema Linking
Listing LABEL:lst:zero_shot_linking is a schema linking prompt template proposed in C3 [9]. Listing LABEL:lst:few_shot_linking is used for few-shot learning proposed in DIN-SQL [33]. In addition, we conduct tests on the PreSQL method. This method generates an SQL query statement based on the user query and database schema, and extracts the corresponding tables from that statement.
@@ B.2 Additional Experimental Results on BIRD
To further justify the effectiveness and robustness of the core conclusions of this paper, we conduct identical experiments on the BIRD development dataset. The BIRD dev set contains 1,534 instances, which exhibit a larger scale compared to “BigTable-0.2k”. In the experiments, we only choose representative LLMs and methods that are influential in deriving core conclusions. Results for five sub-tasks are shown in Table 10, Figure 6, tables˜11, 12 and 13, respectively. These experimental results are largely consistent with the core conclusions obtained on “BigTable-0.2k”, supporting the validity of the conclusions.
@@ C.2 Details for SQL Debugging
Incorrect
Tables
The tables you inquired about is incorrect.
@@
Incorrect
Columns
You have found the correct tables.But you select wrong columns.
@@ C.4 Details for SQL-to-Text
  • Rouge [25] focuses on token-level syntactical similarity by breaking down the predicted question and the GT question into N-grams and calculating the recall rate. We use ROUGE-1, ROUGE-2, and ROUGE-L, respectively.
  • BertScore [59] focuses on semantic similarity between predicted and GT questions. Based on the BERT pre-trained model, BertScore uses contextual embeddings to describe sentences and calculates the cosine similarity between them.

### v2

@@ 4.2 SQL Debugging
In recent studies, researchers have demonstrated that LLMs possess self-reflection and self-correction capabilities similar to those of humans Yao et al. (2022); Shinn et al. (2023); Zhang et al. (2023a). Additionally, previous studies also investigate the potential of LLMs to debug the code they generate Chen et al. (2023); Pourreza and Rafiei (2024). In this section, we provide a comprehensive analysis of the performance of numerous SQL debugging methods across LLMs.
@@ 4.2.2 Debug Evaluation
Dataset normal-$\longrightarrow$ SQLCoder-34B Codellama-34B InternLM-70B Llama2-Chat-70B
@@
Regenerate 0.00 12.42 13.21 18.75
w/ Wrong SQL 0.66 2.48 6.92 7.95
w/ Wrong SQL + System_error_info 4.64 11.18 9.43 11.36
w/ Wrong SQL + All_error_info 4.64 11.80 8.81 10.80
@@
Regenerate 9.27 11.80 0.63 15.34
w/ Wrong SQL 7.95 10.56 5.66 9.66
w/ Wrong SQL + System_error_info 10.60 10.56 8.81 11.93
w/ Wrong SQL + All_error_info 11.92 11.80 8.81 13.07
@@ 4.3 SQL Optimization
Execution efficiency of SQL queries is a critical aspect, particularly in real-time systems that utilize large-scale databases. In this section, we further explore whether LLMs are able to enhance the execution efficiency of correct SQL queries. Formally, the SQL optimization process Pirahesh et al. (1992) involves transforming the initial SQL queries $\mathcal{Y}$ into an optimized form, denoted as $\mathcal{Y}^{\rm o}$, with the goal of improving efficiency while maintaining identical results:
@@
VES is commonly employed to evaluate the efficiency of SQL query execution. However, in practice, LLMs can sometimes rewrite a correct SQL query into an incorrect one, making it challenging to figure out if the main reason for the decline in VES is due to these incorrect rewrites or a decrease in the SQL execution efficiency. To this end, we suggest adopting a complementary metric C-VES (Correct-VES):
@@
Methods Prompt Template Metrics ChatGPT SQLCoder-34B Codellama-34B InternLM-70B InternLM2-20B
Baseline SimpleDDL-MD-Chat VES 36.90 24.28 18.94 22.63 19.81
C-VES 102.50 101.17 99.68 110.39 94.33
Two-Stage Generation zero-shot with $\mathcal{Y}$ VES 30.73 16.86 17.34 15.33 20.28
C-VES 102.43 102.19 102.08 102.17 101.42
w/ $\mathcal{Y}+\mathcal{S}+\mathcal{Q}$ VES 32.21 18.90 19.28 18.44 20.77
C-VES 102.24 102.14 101.47 102.43 101.30
few-shot w/ demo VES 32.19 18.84 18.34 17.97 20.86
C-VES 102.18 101.84 101.88 102.68 101.75
w/ demo + comments VES 32.65 18.28 18.52 17.45 20.92
C-VES 102.03 101.54 102.86 102.66 102.06
Direct Generation SimpleDDL-MD-Chat-Efficiency VES 39.26 27.77 20.75 25.23 25.93
C-VES 103.31 102.84 103.75 102.98 101.70
@@ 4.4 SQL-to-Text
The goal of SQL-to-Text is to transform the SQL query back into its original natural language question Xu et al. (2018); Ma et al. (2021); Shu et al. (2021). While it seems that SQL-to-Text cannot serve as a sub-task within the Text-to-SQL pipeline to improve the performance of End-to-End Text-to-SQL systems, employing this conversion as a supplementary step within the pipeline can indeed provide valuable insights. By converting the generated SQL statements back into text and juxtaposing these with the semantics of the original user questions, we can assess the accuracy of the SQL statements produced. In addition, it can assist researchers in evaluating the semantic comprehension capabilities of different LLMs, thus facilitating the development of more effective Text-to-SQL methodologies.
@@
ChatGPT SQLCoder-34B Codellama-34B InternLM-70B InternLM2-20B Llama2-Chat-70B
Rouge-1 0.428 0.118 0.359 0.408 0.410 0.368
Rouge-2 0.384 0.039 0.167 0.198 0.294 0.159
Rouge-L 0.545 0.108 0.320 0.372 0.358 0.325
BertScore 0.887 0.779 0.856 0.886 0.887 0.885
ChatGPT Evaluator 70.5% 8.5% 47.0% 63.5% 66.5% 50.0%
InternLM2 Evaluator 62% 22.5% 49.8% 65.8% 71.3% 60.0%
To this end, we assess the performance of SQL-to-Text across different LLMs (See Appendix A.4 for prompt templates). The selected metrics for evaluation encompass the F1 values of Rouge-1/2/L and BertScore, along with the application of LLM to assess the semantic coherence between the two texts. The evaluation results are depicted in Table 7. ChatGPT and InternLM2 demonstrate the highest performance, followed by InternLM, while Codellama and SQLCoder exhibit comparatively lower performance. This highlights that even in regards to semantic description of code, general-purpose models exhibit significantly stronger descriptive capabilities compared to coding-specific models.
@@ 4.5 Schema Linking
Schema linking is recognized as a crucial prerequisite of generating correct SQL queries. It involves aligning entity references in the question with the corresponding schema tables or columns, requiring the model to understand both structure and value of the database, as well as the the semantics of user questions. In LLM-based Text-to-SQL, prior studies Dong et al. (2023); Pourreza and Rafiei (2024) design prompt instructions with in-context learning examples to enable LLMs to retrieve linked tables and columns, which are then used for the downstream Text-to-SQL task. However, none of these methods individually evaluate the performance of schema linking, and explicit evaluation metrics have yet to be established. Moreover, despite considerable advancements in semantic comprehension and generalization brought by LLMs, the performance of schema linking is still far from promising.
@@ 4.5.1 Evaluation Metric for Schema Linking: RES
What schema linking method is considered good? A straightforward goal of schema linking is that the GT tables should be retrieved as much as possible. However, due to the ambiguity inherent in natural language and the potential semantic similarities between candidate tables, retrieving more GT tables typically comes at the cost of a higher redundancy, known as the precision-recall trade-off. As discussed in Pourreza and Rafiei (2024), excessive table retrieval can introduce redundant joins between tables, potentially impairing the EX of Text-to-SQL generation.
@@
where $T_{n}$ and $\hat{T}_{n}$ denote the set of GT tables and retrieved tables for the $n$-th instance, respectively, ||$\cdot$|| ⋅ | refers to the scale of a set.
@@ 4.5.2 Schema Linking Evaluation
  • Zero Shot: A schema linking prompt proposed in C3 Dong et al. (2023), which instructs the LLM to rank all the tables from relevant to irrelevant in a zero-shot manner.
  • Few Shot: Instructing the LLM to retrieve only the most important tables with few-shot demonstration, presented in DIN-SQL Pourreza and Rafiei (2024).
@@
Note that we organize the information of the database schema in a way similar to "SimpleDDL-" in the prompt of all the methods mentioned above, which ignores the information about foreign keys. However, as argued in Wang et al. (2019), foreign keys embody features of known schema relations and provide importance clues for understanding the database structure. To this end, we conduct experiments under both settings, w/ and w/o foreign keys, to investigate how incorporating foreign keys in the prompt influences the performance of schema linking. Results are demonstrated in Table 8 (refer to Appendix B for results on detailed metrics like Exact Match & Subset Match).
@@
Can foreign key information facilitate schema linking? The introduction of foreign key information yield improved performance across all methods and all LLMs. This is evident since a valid JOIN operation in SQL queries is typically based on foreign keys. The foreign key information helps the model retrieve more ground truth tables by indicating all potential table pairs involved in a JOIN operation.
@@ A.4 SQL-to-Text
Listing LABEL:lst:sql_to_text is the prompt template used for SQL-to-Text task. Listing LABEL:lst:Semantic_consistency is utilized to assess the consistency between the transformed text description and the original user query semantics. To mitigate the influence of the order of the two text inputs, we employ ChatGPT and InternLM2 to conduct two sequential tests and calculate the average of the results.
@@ A.5 Schema Linking
Listing LABEL:lst:zero_shot_linking is a schema linking prompt template proposed in C3 Dong et al. (2023). Listing LABEL:lst:few_shot_linking is used for few-shot learning proposed in DIN-SQL Pourreza and Rafiei (2024). In addition, we conduct tests on the PreSQL method. This method generates an SQL query statement based on the user query and database schema, and extracts the corresponding tables from that statement.
@@ B.2 Additional Experimental Results on BIRD
To further justify the effectiveness and robustness of the core conclusions of this paper, we conduct identical experiments on the BIRD development dataset. The BIRD dev set contains 1,534 instances, which exhibit a larger scale compared to “BigTable-0.2k”. In the experiments, we only choose representative LLMs and methods that are influential in deriving core conclusions. Results for five sub-tasks are shown in Table 10, Figure 6, tables 11, 12 and 13, respectively. These experimental results are largely consistent with the core conclusions obtained on “BigTable-0.2k”, supporting the validity of the conclusions.
@@ C.2 Details for SQL Debugging
Incorrect Tables The tables you inquired about is incorrect.
@@
Incorrect Columns You have found the correct tables.But you select wrong columns.
@@ C.4 Details for SQL-to-Text
  • Rouge Lin (2004) focuses on token-level syntactical similarity by breaking down the predicted question and the GT question into N-grams and calculating the recall rate. We use ROUGE-1, ROUGE-2, and ROUGE-L, respectively.
  • BertScore Zhang et al. (2019) focuses on semantic similarity between predicted and GT questions. Based on the BERT pre-trained model, BertScore uses contextual embeddings to describe sentences and calculates the cosine similarity between them.